OpenStudy (anonymous):

Can someone check if my answer is correct? Given the linear equation x – y = 5, find the slope of its graph. A. 1 B. 5 C. –1 D. –5 My answer is C. -1 @ccswims

OpenStudy (ccswims):

It's A

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Shouldn't the slope be negative?

OpenStudy (ccswims):

You can't have -y = -x + 5 ... the 'y' can't be negative, so you must divide everything by -1 y = -x(-1) + (-1) y = x - 5

OpenStudy (ccswims):

I meant + 5(-1)

OpenStudy (anonymous):

Oh! Okay! Thanks again!